OpenStudy (anonymous):

Find the limit of the function by using direct substitution. lim-> pi/2 (2e^x cosx) 0 1 2e^(pi/2) pi/2

OpenStudy (welshfella):

2 * e^(pi/2) * cos (pi/2) make sure to set your calculator to radians

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \lim_{x\rightarrow \pi/2}~2e^x\cos(x) =2e^{\pi/2}\cos(\pi/2) }\)

OpenStudy (astrophysics):

What does cos(pi/2) give you?

OpenStudy (astrophysics):

That should be a give away

OpenStudy (solomonzelman):

This is what direct substitution means here. When you have a limit \(\large\color{black}{ \displaystyle \lim_{x\rightarrow a}~f(x) }\) then, by direct substitution, your limit is equal to \(\large f(a)\).

OpenStudy (solomonzelman):

(most limits as you have previously seen or going to see, you can't apply direct substitution tho', whether that is fortunate or not....)

OpenStudy (solomonzelman):

(( We can do a taylor approximation here for \(\pi/2\) and for e, if you would like ))

OpenStudy (solomonzelman):

I will tell you this. If you have: \(\large\color{black}{ \displaystyle \lim_{x\rightarrow a}~f(x) }\) then, for any continous function f(x), direct substitution will work. (unless you have a=±∞)

OpenStudy (anonymous):

i got c is that correct? @SolomonZelman @Astrophysics

OpenStudy (solomonzelman):

\(\cos(90)=?\)

OpenStudy (solomonzelman):

if you don't remember, then, unit circle, or: \(\color{black}{ \displaystyle \cos(90)=\cos(45+45)=\cos(45)\cos(45)-\sin(45)\sin(45)=0 }\) (knowing that `sin(45)=cos(45)` => = √2/2)

OpenStudy (anonymous):

1?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \lim_{x\rightarrow \pi/2}~2e^x\cos(x) =2e^{\pi/2}\cdot \cos(\pi/2) =2e^{\pi/2}\cdot 0=?}\)

OpenStudy (anonymous):

0

OpenStudy (solomonzelman):

yes, this limit is equivalent to zero.

OpenStudy (anonymous):

thank u!!

OpenStudy (solomonzelman):

(yes, 0, from both sides.)

OpenStudy (solomonzelman):

Welcome into the calculus world:) xD Good luck with your math!